### v4

@@ A.2 Proofs in Section 4
Proof [Proof of Lemma 4.2] Following the classical dynamical programming principle argument in, for instance, oksendal2007applied , we have the following the HJB equation associated with (4.9), where for any $x\in\mathbb{R}$,
@@
Proof [Proof of Theorem 4.3] For any $(\pi,\bm{\mu})\in\mathcal{A}$, applying Itô’s formula (e.g., (oksendal2007applied, , Theorem 1.14.)) to $g(t,X_{t})$ subject to (4.6), with $g(t,y)=e^{-rt}V(y)$ for any $(t,y)\in[0,+\infty)\times\mathbb{R}$, we have
@@
$\hskip 10.0pt-\int_{0}^{t}e^{-rs}\left\{f(X_{s-}-\lambda_{1}\mathcal{R}(\pi_{s})+\pi_{s}\left[\mathbb{E}_{\xi\sim\mu_{s}}\left[l(\xi)\right]-\lambda_{2}\mathcal{H}(\mu_{s};X_{s-})\right])\right\}\mathrm{d}s.$
Since $\{\int_{0}^{t}e^{-rs}V^{\prime}(X_{s})\sigma(X_{s-})\mathrm{d}W_{s}\}_{t\geq 0}$, taking expectation on both sides of the inequality, we have
@@
$V(x)\leq J^{\bm{\lambda}}(x;\pi,\bm{\mu}),\ \ \forall x\in\mathbb{R}\implies V\leq\hat{\psi}^{\lambda}.$
@@
$V(x)=J^{\bm{\lambda}}(x;\pi^{*},\bm{\mu}^{*}),\ \ \forall x\in\mathbb{R}\implies V=\hat{\psi}^{\bm{\lambda}}.$
@@ A.3 Proofs in Section 5
Therefore, $\psi^{\bm{\lambda},0}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Suppose by induction that $\psi^{\bm{\lambda},n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$ for $n\in\mathbb{N}^{+}$. By Theorem 3.5 and Assumption 2.3-(iii), $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Furthermore, one can conclude that $\psi^{\bm{\lambda},n}=T^{\lambda_{1}}[\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}]\in\mathscr{L}_{\frac{L_{f}}{r-G}}$ based on Lemma 3.4, which closes the proof.
@@
Proof [Proof of Theorem 5.3] For $\hat{\psi}^{\bm{\lambda}}$, 0 is a natural lower bound. From Theorem 5.2, we know that $\hat{\psi}^{\bm{\lambda}}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. By Theorem 3.5, we know that for each $x\in\mathbb{R}$, $\mathcal{M}^{\lambda_{2}}\hat{\psi}^{\bm{\lambda}}$ is attained at and only at $\mu^{*}_{x}$.
@@ A.4 Proofs in Section 6
Inductively suppose that $\psi^{n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Then by (guo2009smooth, , Lemma 2.2-(3)), $\mathcal{M}\psi^{n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Subsequently, using Lemma 2.2, the following holds for any $\tau$:
@@
$\leq$ $\mathbb{E}\left[\int_{0}^{\tau}e^{-rs}L_{f}|X_{s}-X^{\prime}_{s}|\mathrm{d}s+e^{-r\tau}L_{\psi^{n-1}}|X_{\tau}-X^{\prime}_{\tau}|\right]$
$\leq$ $\mathbb{E}\left[\sup_{s}e^{-Cs}|X_{s}-X_{s}^{\prime}|\right]\cdot\left(\int_{0}^{\infty}e^{-(r-G)s}L_{f}\mathrm{d}s+e^{-(r-G)\tau}\frac{L_{f}}{r-G}\right)$
$\leq$ $\frac{L_{f}}{r-G}|x-x^{\prime}|.$
@@
Let us take the constant suboptimal control $\bar{\pi}=\frac{1}{\lambda_{1}^{2}}$, we have
@@
$\leq$ $\;\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})}\left(f(X_{t})+\frac{1}{\lambda_{1}^{2}}\mathcal{M}^{\lambda_{2}}\psi^{n-1}(X_{t})-\lambda_{1}\mathcal{R}\left(\frac{1}{\lambda_{1}^{2}}\right)\right)\mathrm{d}t\right]$
$\leq$ $\;\mathbb{E}_{x}\bigg[\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})}\bigg(\frac{1}{\lambda_{1}^{2}}\left(\mathcal{M}^{\lambda_{2}}\psi^{n-1}(x)+\mathcal{M}^{\lambda_{2}}\psi^{n-1}(X_{t})-\mathcal{M}^{\lambda_{2}}\psi^{n-1}(x)\right)+M\bigg)\mathrm{d}t\bigg]$
@@
$\leq$ $\;\mathcal{M}^{\lambda_{2}}(x)+M\lambda_{1}^{2}+\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})t}\left(\frac{1}{\lambda_{1}^{2}}\left(\mathcal{M}^{\lambda_{2}}\psi^{n-1}(X_{t})-\mathcal{M}^{\lambda_{2}}\psi^{n-1}(x)\right)\right)\mathrm{d}t\right]$
@@
$\leq$ $\;\mathcal{M}^{\lambda_{2}}\psi^{n}(x)+M\lambda_{1}^{2}+\frac{2L_{f}M}{r-G}\lambda_{1}+C\lambda_{1}\log\frac{1}{\lambda_{1}},$
where the second inequality comes from $\|f\|_{L^{\infty}(\mathbb{R})}\leq M$ by Assumption 6.1, the third inequality is due to $\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})t}\mathrm{d}t=\frac{\lambda_{1}^{2}}{1+r\lambda_{1}^{2}}\leq C\lambda_{1}^{2}$, and the fourth inequality comes from (A.10). For sufficiently small $\lambda_{1}>0$, we have $\mathcal{T}^{\bm{\lambda}}\psi^{n}(x)\leq\mathcal{M}^{\lambda_{2}}\psi^{n}(x)+C\lambda_{1}\log\frac{1}{\lambda_{1}}.$
Proof [Proof of Proposition 6.8] Recall that $\mathcal{T}^{\bm{\lambda}}\psi^{n}$ and $\tilde{\psi}^{n}$ respectively solves (5.3) and (6.4) with obstacle function $\mathcal{M}^{\lambda_{2}}\psi^{n}$.
@@
When $n=0$, it holds that $\psi^{\bm{\lambda},0}\equiv\psi^{0}$. Suppose $\psi^{\bm{\lambda},n}\geq\psi^{n}-\frac{\lambda_{1}}{r}$. Set $w^{n}=\psi^{n}-\frac{\lambda_{1}}{r}$. By the value iteration (6.2), we have
$\mathcal{M}^{\lambda_{2}}w^{n}=\mathcal{M}^{\lambda_{2}}\left(\psi^{n}-\frac{\lambda_{1}}{r}\right)=\mathcal{M}^{\lambda_{2}}\psi^{n}-\frac{\lambda_{1}}{r}\geq\psi^{n+1}-\frac{\lambda_{1}}{r}=w^{n+1},$
@@
Put $\alpha:=\lambda_{1}\log(1/\lambda_{1})$ and let $\widetilde{\mu}:=\max\{1-\mu,\tfrac{1}{2}\}<1$. Both exponential terms on the RHS of (A.16) satisfy $2^{-n(\lambda_{1})},(1-\mu)^{n(\lambda_{1})}\leq\widetilde{\mu}^{\,n(\lambda_{1})}$. Applying (reisinger2020error, , Lemma 4.9) with $\gamma=1$ yields a constant $C_{0}$ such that
@@ A.5 Proof in Section 7
Proof [Proof of Theorem 7.1] Denote $w^{n}(x)=\psi^{\bm{\lambda},n+1}(x)-\psi^{\bm{\lambda},n}(x)$, then for any $x\in\mathbb{R}$,
$\ \mathcal{L}w^{n}(x)-rw^{n}(x)$ $=\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\psi^{\bm{\lambda},n+1}(x)}{\lambda_{1}}\right)-\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)-\psi^{\bm{\lambda},n}(x)}{\lambda_{1}}\right)$
$=\exp\left(-\frac{c}{\lambda_{1}}\right)\left(\psi^{\bm{\lambda},n+1}(x)-\psi^{\bm{\lambda},n}(x)-(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x))\right),$
where the second equality holds by mean value theorem, with $c$ lying between $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\psi^{\bm{\lambda},n+1}(x)$ and $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)-\psi^{\bm{\lambda},n}(x)$.
Note that both $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)-\psi^{\bm{\lambda},n}(x)$ and $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\psi^{\bm{\lambda},n+1}(x)$ are no more than $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(x)-\psi^{\bm{\lambda},0}(x)$, we further argue that $\exp(-(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(x)-\psi^{\bm{\lambda},0}(x))/\lambda_{1})\coloneqq C_{0}$ is bounded, which will be invoked later.
Let us derive the lower bound of $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}$:
$\mathcal{M}^{{\lambda}_{2}}\psi^{\bm{\lambda}}(x)=-\lambda_{2}\log\mathbb{E}_{\xi\sim\Phi_{x}}\!\left[\exp\!\left(-\frac{\psi^{\bm{\lambda}}(x+\xi)+l(\xi)}{\lambda_{2}}\right)\right]$
$=$ $\;\psi^{\bm{\lambda}}(0)+l(-x)-\lambda_{2}\log\mathbb{E}_{\xi\sim\Phi_{x}}\!\left[\exp\!\left(-\frac{\psi^{\bm{\lambda}}(x+\xi)+l(\xi)-\psi^{\bm{\lambda}}(0)-l(-x)}{\lambda_{2}}\right)\right]$
As $\psi^{\bm{\lambda}}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$,
$\exp\!\left(-\frac{\psi^{\bm{\lambda}}(x+\xi)+l(\xi)-\psi^{\bm{\lambda}}(0)-l(-x)}{\lambda_{2}}\right)\leq\exp\left(\frac{\left(\frac{L_{f}}{r-G}+L_{l}\right)|x+\xi|}{\lambda_{2}}\right).$
As $x+\xi\sim\Phi_{0}$, one can derive that its expectation is bounded. This gives the lower boundedness of $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(x)$, which leads to a bound on $C_{0}$. Namely,
@@
Since $w^{n}$ is bounded, we may assume that $w^{n}(x)$ attains its maximum at $\overline{x}$, then by first- and second-order optimality conditions
$0$ $\geq\mathcal{L}w^{n}(\overline{x})$
@@
$=\left(r+\exp\left(-\frac{c}{\lambda_{1}}\right)\right)w^{n}(\overline{x})-\exp\left(-\frac{c}{\lambda_{1}}\right)(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(\overline{x})-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(\overline{x})).$
Rearranging the above inequality,

### v5

@@ A.2 Proofs in Section 4
Proof [Proof of Lemma 4.2] Following the classical dynamical programming principle argument in, for instance, Øksendal and Sulem (2007), we have the following the HJB equation associated with (4.9), where for any $x\in\mathbb{R}$,
@@
Proof [Proof of Theorem 4.3] For any $(\pi,\bm{\mu})\in\mathcal{A}$, applying Itô’s formula (e.g., (Øksendal and Sulem, 2007, Theorem 1.14.)) to $g(t,X_{t})$ subject to (4.6), with $g(t,y)=e^{-rt}V(y)$ for any $(t,y)\in[0,+\infty)\times\mathbb{R}$, we have
@@
$\hskip 10.0pt-\int_{0}^{t}e^{-rs}\left\{f(X_{s-})-\lambda_{1}\mathcal{R}(\pi_{s})+\pi_{s}\left[\mathbb{E}_{\xi\sim\mu_{s}}\left[l(\xi)\right]-\lambda_{2}\mathcal{H}(\mu_{s};X_{s-})\right])\right\}\mathrm{d}s.$
Since $\{\int_{0}^{t}e^{-rs}V^{\prime}(X_{s})\sigma(X_{s-})\mathrm{d}W_{s}\}_{t\geq 0}$ is a martingale, taking expectation on both sides of the inequality, we have
@@
$V(x)\leq J^{\bm{\lambda}}(x;\pi,\bm{\mu}),\ \ \forall x\in\mathbb{R}\implies V\leq\tilde{\psi}^{\lambda}.$
@@
$V(x)=J^{\bm{\lambda}}(x;\pi^{*},\bm{\mu}^{*}),\ \ \forall x\in\mathbb{R}\implies V=\tilde{\psi}^{\bm{\lambda}}.$
@@ A.3 Proofs in Section 5
Therefore, $\psi^{\bm{\lambda},0}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Suppose by induction that $\psi^{\bm{\lambda},n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$ for $n\in\mathbb{N}^{+}$. By Lemma 3.2 and Assumption 2.3-(iii), $L_{l}\leq\frac{L_{f}}{r-G}$, and hence $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Furthermore, one can conclude that $\psi^{\bm{\lambda},n}=T^{\lambda_{1}}[\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}]\in\mathscr{L}_{\frac{L_{f}}{r-G}}$ based on Lemma 3.4, which closes the proof.
@@
Proof [Proof of Theorem 5.3] For $\hat{\psi}^{\bm{\lambda}}$, 0 is a natural lower bound. From Theorem 5.2, we know that $\hat{\psi}^{\bm{\lambda}}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Following a similar line of proof as Lemma 3.2, we know that for each $x\in\mathbb{R}$, $\mathcal{M}^{\lambda_{2}}\hat{\psi}^{\bm{\lambda}}$ is attained at and only at $\mu^{*}_{x}$.
@@ A.4 Proofs in Section 6
Inductively, suppose that $\psi^{n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Define the cost functional for optimal stopping problem at step $n$ corresponding to (6.1): For any $x\in\mathbb{R}$, $\tau$ as $\mathbb{F}$-stopping time,
$J^{n}(x;\tau):=\mathbb{E}_{x}\left[\int_{0}^{\tau}e^{-rt}f(X_{t})\mathrm{d}t+e^{-r\tau}\mathcal{M}\psi^{n-1}(X_{\tau})\right].$
Then by (Guo and Wu, 2009, Lemma 2.2-(3)), $\mathcal{M}\psi^{n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Subsequently, using Lemma 2.2, the following holds for any $\tau$:
@@
$\leq$ ${\mathbb{E}\left[\sup_{s\geq 0}e^{-Gs}|X_{s}-X_{s}^{\prime}|\right]\cdot\left(\int_{0}^{\tau}e^{-(r-G)s}L_{f}\mathrm{d}s+e^{-(r-G)\tau}\frac{L_{f}}{r-G}\right)\leq\frac{L_{f}}{r-G}|x-x^{\prime}|}.$
@@
Recall that $\mathcal{T}^{\bm{\lambda}}\psi^{n}$, by the definition of the randomized compound optimal stopping operator $\mathcal{T}^{\bm{\lambda}}$, solves the problem
$\mathcal{T}^{\bm{\lambda}}\psi^{n}(x)=\inf_{\pi}\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-rt}p_{t}\left(f(X_{t})-\lambda_{1}\mathcal{R}(\pi_{t})+\pi_{t}\mathcal{M}^{\lambda_{2}}\psi^{n}(X_{t})\right)\mathrm{d}t\right],\quad\forall x\in\mathbb{R},$
subject to (2.1) and (2.12) with $p_{0}=1$. Moreover, $\mathcal{T}^{\bm{\lambda}}\psi^{n}$ is the unique $\mathcal{C}^{2}$ solution to the HJB equation
$(\mathcal{L}-r)u+f-\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}\psi^{n}-u}{\lambda_{1}}\right)=0.$
Taking the constant suboptimal control $\bar{\pi}=\frac{1}{\lambda_{1}^{2}}$ for the problem $\mathcal{T}^{\bm{\lambda}}\psi^{n}(x)$ yields
@@
$\leq$ $\;\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})}\left(f(X_{t})+\frac{1}{\lambda_{1}^{2}}\mathcal{M}^{\lambda_{2}}\psi^{n}(X_{t})-\lambda_{1}\mathcal{R}\left(\frac{1}{\lambda_{1}^{2}}\right)\right)\mathrm{d}t\right]$
$\leq$ $\;\mathbb{E}_{x}\bigg[\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})}\bigg(\frac{1}{\lambda_{1}^{2}}\left(\mathcal{M}^{\lambda_{2}}\psi^{n}(x)+\mathcal{M}^{\lambda_{2}}\psi^{n}(X_{t})-\mathcal{M}^{\lambda_{2}}\psi^{n}(x)\right)+M\bigg)\mathrm{d}t\bigg]$
@@
$\leq$ $\;\mathcal{M}^{\lambda_{2}}\psi^{n}(x)+M\lambda_{1}^{2}+\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})t}\left(\frac{1}{\lambda_{1}^{2}}\left(\mathcal{M}^{\lambda_{2}}\psi^{n}(X_{t})-\mathcal{M}^{\lambda_{2}}\psi^{n}(x)\right)\right)\mathrm{d}t\right]$
@@
$\leq$ $\;\mathcal{M}^{\lambda_{2}}\psi^{n}(x)+M\lambda_{1}^{2}+\frac{CL_{f}M}{r-G}\lambda_{1}+C\lambda_{1}\log\frac{1}{\lambda_{1}},$
where the second inequality comes from $\|f\|_{L^{\infty}(\mathbb{R})}\leq M$ by Assumption 6.1, the third inequality is due to $\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})t}\mathrm{d}t=\frac{\lambda_{1}^{2}}{1+r\lambda_{1}^{2}}\leq C\lambda_{1}^{2}$, and the fourth inequality can be derived by:
$\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})t}\frac{1}{\lambda_{1}^{2}}\left(\mathcal{M}^{\lambda_{2}}\psi^{n}(X_{t})-\mathcal{M}^{\lambda_{2}}\psi^{n}(x)\right)\mathrm{d}t\right]$
$\leq$ $\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})t}\frac{1}{\lambda_{1}^{2}}\frac{L_{f}M}{r-G}(t+\sqrt{t})\mathrm{d}t$
$=$ $\frac{1}{\lambda_{1}^{2}}\frac{L_{f}M}{r-G}\left(\frac{1}{(r+\lambda_{1}^{-2})^{2}}+\frac{\sqrt{\pi}}{(r+\lambda_{1}^{-2})^{3/2}}\right)$
$\leq$ $\frac{CL_{f}M}{r-G}\lambda_{1}$
for some constant $C>0$.
Finally, it can be concluded from (A.11) that for sufficiently small $\lambda_{1}>0$, we have $\mathcal{T}^{\bm{\lambda}}\psi^{n}(x)\leq\mathcal{M}^{\lambda_{2}}\psi^{n}(x)+C\lambda_{1}\log\frac{1}{\lambda_{1}}.$
Proof [Proof of Proposition 6.8] Recall that $\mathcal{T}^{\bm{\lambda}}\psi^{n}$ solves (5.3) with obstacle function $\mathcal{M}^{\lambda_{2}}\psi^{n}$:
$(\mathcal{L}-r)\mathcal{T}^{\bm{\lambda}}\psi^{n}+f-\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}\psi^{n}-\mathcal{T}^{\bm{\lambda}}\psi^{n}}{\lambda_{1}}\right)=0.$
Meanwhile, $\tilde{\psi}^{n}$ solves (6.4):
$\min\{\mathcal{L}\tilde{\psi}^{n}-r\tilde{\psi}^{n}+f,\mathcal{M}^{\lambda_{2}}\psi^{n}-\tilde{\psi}^{n}\}=0.$
@@
When $n=0$, it holds that $\psi^{\bm{\lambda},0}\equiv\psi^{0}$. Suppose $\psi^{\bm{\lambda},n}\geq\psi^{n}-\frac{\lambda_{1}}{r}$. Set $w^{n}=\psi^{n}-\frac{\lambda_{1}}{r}$. By the value iteration (6.2) and Lemma 3.3-(ii), we have
$\mathcal{M}^{\lambda_{2}}w^{n}=\mathcal{M}^{\lambda_{2}}\left(\psi^{n}-\frac{\lambda_{1}}{r}\right)=\mathcal{M}^{\lambda_{2}}\psi^{n}-\frac{\lambda_{1}}{r}\geq\mathcal{M}\psi^{n}-\frac{\lambda_{1}}{r}\geq\psi^{n+1}-\frac{\lambda_{1}}{r}=w^{n+1},$
@@
Put $\alpha:=\lambda_{1}\log(1/\lambda_{1})$ and let $\widetilde{\mu}:=\max\{1-\mu,\tfrac{1}{2}\}<1$. Both exponential terms on the RHS of (A.16) satisfy $2^{-n(\lambda_{1})},(1-\mu)^{n(\lambda_{1})}\leq\widetilde{\mu}^{\,n(\lambda_{1})}$. Applying (Reisinger and Zhang, 2020, Lemma 4.9) with $\gamma=1$ yields a constant $C_{0}$ such that
@@ A.5 Proof in Section 7
Proof [Proof of Theorem 7.1] Denote $w^{n}(x)=\psi^{\bm{\lambda},n+1}(x)-\psi^{\bm{\lambda},n}(x)$, $g^{n}(x)=\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)$. Fix $n\in\mathbb{N}^{+}$, for any $x\in\mathbb{R}$,
$\mathcal{L}w^{n}(x)-rw^{n}(x)$
$=$ $\;\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\psi^{\bm{\lambda},n+1}(x)}{\lambda_{1}}\right)-\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)-\psi^{\bm{\lambda},n}(x)}{\lambda_{1}}\right)$
$=$ $\;\exp\left(-\frac{c(x)}{\lambda_{1}}\right)\left(\psi^{\bm{\lambda},n+1}(x)-\psi^{\bm{\lambda},n}(x)-(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x))\right),$
where the second equality holds by mean value theorem, with $c(x)$ pointwisely lying between $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\psi^{\bm{\lambda},n+1}(x)$ and $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)-\psi^{\bm{\lambda},n}(x)$.
By Lemma 3.1-(iv) and 5.2-(i), both $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)-\psi^{\bm{\lambda},n}(x)$ and $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\psi^{\bm{\lambda},n+1}(x)$ are pointwisely no less than $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(x)-\psi^{\bm{\lambda},0}(x)$. Therefore, $c(x)\geq\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(x)-\psi^{\bm{\lambda},0}(x)$. We further find a uniform lower bound for $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(x)-\psi^{\bm{\lambda},0}(x)$ to argue that $C_{0}(x)=:\exp(-(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(x)-\psi^{\bm{\lambda},0}(x))/\lambda_{1})$ is uniformly upper bounded, which will be invoked later.
To derive the lower bound of $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}$, notice that it is equivalent to
$\psi^{\bm{\lambda}}(0)+l(-x)-\lambda_{2}\log\mathbb{E}_{\xi\sim\Phi_{x}}\!\left[\exp\!\left(-\frac{\psi^{\bm{\lambda}}(x+\xi)+l(\xi)-\psi^{\bm{\lambda}}(0)-l(-x)}{\lambda_{2}}\right)\right].$
As $\psi^{\bm{\lambda}}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$, we have $\exp\!\left(-\frac{\psi^{\bm{\lambda}}(x+\xi)+l(\xi)-\psi^{\bm{\lambda}}(0)-l(-x)}{\lambda_{2}}\right)\leq\exp\left(\frac{\left(\frac{L_{f}}{r-G}+L_{l}\right)|x+\xi|}{\lambda_{2}}\right)$. As $x+\xi\sim\Phi_{0}$, one can derive that its expectation is bounded. This gives the uniform lower boundedness of $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}$. Combining with the upper bound on $\psi^{\bm{\lambda},0}$ by proposition 6.5 leads to a uniform upper bound on $C_{0}(x)$, denoted as $C_{0}$. Namely,
@@
In the following, we prove that
$\|w^{n}\|_{L^{\infty}(\mathbb{R})}\leq\frac{C_{0}}{C_{0}+r}\,\|\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}\|_{L^{\infty}(\mathbb{R})}=:K_{n}.$
Let us first show $\sup_{x\in\mathbb{R}}w^{n}(x)\leq K_{n}$. Since $w^{n}$ is bounded, for each $\epsilon>0$ the function $v^{n,\epsilon}(x):=w^{n}(x)-\frac{\epsilon}{2}x^{2}$ satisfies $v^{n,\epsilon}(x)\to$-$\infty$ as $|x|\to\infty$ and hence attains its maximum at some point $\overline{x}_{\epsilon}\in\mathbb{R}$: $v^{n,\epsilon}(\overline{x}_{\epsilon})=\max_{x\in\mathbb{R}}v^{n,\epsilon}(x).$
Recall that
$\mathcal{L}w^{n}(x)-rw^{n}(x)=a^{n}(x)\bigl(w^{n}(x)-g^{n}(x)\bigr),$
where
$a^{n}(x):=\exp\!\Big(-\tfrac{c(x)}{\lambda_{1}}\Big)\in(0,C_{0}],\qquad g^{n}(x):=\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x),$
and set $B_{n}:=\|g^{n}\|_{L^{\infty}(\mathbb{R})}$ so that $K_{n}=\tfrac{C_{0}}{r+C_{0}}B_{n}$. At the maximum point $\overline{x}_{\epsilon}$ we have
$(v^{n,\epsilon})^{\prime}(\overline{x}_{\epsilon})=0,\quad(v^{n,\epsilon})^{\prime\prime}(\overline{x}_{\epsilon})\leq 0,\quad(w^{n})^{\prime}(\overline{x}_{\epsilon})=\epsilon\,\overline{x}_{\epsilon},\quad(w^{n})^{\prime\prime}(\overline{x}_{\epsilon})\leq\epsilon.$
By Assumption 6.1,
$\mathcal{L}w^{n}(\overline{x}_{\epsilon})=\frac{1}{2}\sigma^{2}(\overline{x}_{\epsilon})(w^{n})^{\prime\prime}(\overline{x}_{\epsilon})+b(\overline{x}_{\epsilon})(w^{n})^{\prime}(\overline{x}_{\epsilon})\leq\frac{1}{2}\bar{\sigma}^{2}\,\epsilon+\bar{b}\,\epsilon|\overline{x}_{\epsilon}|\leq C_{1}\,\epsilon\,(1+|\overline{x}_{\epsilon}|),$
for some constant $C_{1}>0$ independent of $n$ and $\epsilon$. Evaluating the identity for $w^{n}$ at $x=\overline{x}_{\epsilon}$ and combining with the above estimate yields
@@
$rw^{n}(\overline{x}_{\epsilon})+\exp\left(-\frac{c}{\lambda_{1}}\right)\bigl(w^{n}(\overline{x}_{\epsilon})-g^{n}(\overline{x}_{\epsilon})\bigr)=\mathcal{L}w^{n}(\overline{x}_{\epsilon})\leq C_{1}\,\epsilon\,(1+|\overline{x}_{\epsilon}|).$
Hence,